# The First Photometric Evidence of a Transient/Variable Source at 5 with JWST

### v1

@@ Abstract
The James Webb Space Telescope (JWST) discovered 79 transients out to 4.8 through the JADES Transient Survey (JTS), but the JTS did not find any 5 transients. Here, we present the first photometric evidence of a 5 transient/variable source with JWST. The source, AT 2023adya, resides in a 5.274 galaxy in GOODS-N, which dimmed from 26.050.02 mag to 26.240.02 mag in the rest-frame optical over approximately two rest-frame months, producing a clear residual signal in the difference image ( 28.010.17 mag; SNvar 6.09) at the galaxy center. Shorter-wavelength bands (F090W/F115W) show no rest-frame ultraviolet brightness change. Based on its rest-frame V-band absolute magnitude of MV 18.48 mag, AT 2023adya could be any core-collapse supernova (SN) subtype or an SN Ia. However, due to low SN Ia rates at high redshift, the SN Ia scenario is unlikely. Alternatively, AT 2023adya may be a variable active galactic nucleus (AGN). However, the JWST NIRCam/Grism spectrum shows no broad H emission line (FWHM 13026 km s-1), disfavoring the variable AGN scenario. It is also unlikely that AT 2023adya is a tidal disruption event (TDE) because the TDE models matching the observed brightness changes have low event rates. Although it is not possible to determine AT 2023adya’s nature based on the two-epoch single-band photometry alone, this discovery indicates that JWST can push the frontier of transient/variable science past 5 and towards the epoch of reionization.
@@ 1 Introduction
Building a sample of high-redshift SNe Ia can provide useful constraints on SN Ia systematics (Riess & Livio, 2006), including whether they are truly similar to low-redshift SNe Ia and can act as standard candles (Pierel et al., 2024a, 2025). This has implications especially for cosmology constraints achieved at higher redshifts, such as the possibility of an evolving dark energy equation of state parameter, w (Adame et al., 2025). CCSNe, on the other hand, closely trace the instantaneous star formation rate and thus provide an independent view of the cosmic star formation rate density. Observed CCSN rates can be compared to those expected from the cosmic star formation rate density to provide constraints on missing CCSN fractions, the progenitor mass range of CCSNe, and the potential scenario of an evolving IMF (Dahlen et al., 2012; Strolger et al., 2015). Additionally, there is evidence that CCSN explosion energies evolve with redshift (Coulter et al., 2025; Moriya et al., 2025). Building a larger sample of high-redshift CCSNe will generate better constraints on how high-redshift CCSNe differ from local CCSNe.
@@
The JWST Advanced Deep Extragalactic Survey (JADES) program obtained two epochs of deep (30 mag) JWST near-infrared camera (NIRCam) images of the Great Observatories Origins Deep Survey South (GOODS-S) field, separated by one observer-frame year (Eisenstein et al., 2023). The JADES Transient Survey (JTS) discovered 38 supernovae (SNe) at 2, and the highest-redshift SN in the JTS sample is associated with a 4.82 host (DeCoursey et al., 2025). This sample includes an SN Ia at 2.9 (Pierel et al., 2024a), a Type Ic broad line (Ic-BL) SN at 2.83 (Siebert et al., 2024), and a likely Type IIP SN at 3.61 (Coulter et al., 2025). Moriya et al. (2025) explores the properties of the robustly-classified Type II SNe in the JTS sample, three of which are at 2. Additionally, within the COSMOS-Web SN survey, an SN Ia has been discovered at 2.15 (Pierel et al., 2025). There have also been multiple serendipitous SN discoveries with JWST (e.g. Yan et al., 2023), including multiply-lensed SNe Ia at 1.78 (Frye et al., 2024; Pierel et al., 2024b; Pascale et al., 2025) and 1.95 (Pierel et al., 2024c).
@@
We conducted a systematic transient and variable search in the GOODS-North field with two sets of overlapping F090W, F115W, and F356W JWST/NIRCam imaging, separated by one observer-frame year (DeCoursey et al., 2024). This search yielded a sample of 39 transient and variable sources, one of which is AT 2023adya. AT 2023adya is associated with a 5.274 host galaxy (Sun et al., 2024), making it the highest-redshift transient/variable source showing a photometric change in brightness discovered thus far with JWST. We explore the types of transient or variable sources that may explain AT 2023adya, which include CCSNe, SNe Ia, variable AGN, and tidal disruption events (TDEs).
@@
In Section 2, we describe the NIRCam observations covering AT 2023adya. Section 3 explains how we discovered AT 2023adya and presents the photometric measurements of AT 2023adya. We describe AT 2023adya’s host and step through the possible sources of AT 2023adya’s transient or variable nature in Section 4, and we summarize our findings in Section 5. Throughout this paper, we express magnitudes using the AB system (Oke & Gunn, 1983) and adopt a flat CDM cosmology with the following parameters: H0 70 km s-1 Mpc-1, 0.7, and 0.3.
@@ 2.1 NIRCam Imaging
The JADES program (program ID: 1181; PI: Eisenstein) obtained 9-band JWST/NIRCam imaging of GOODS-N on UT 2023 February 3, using 4 short-wavelength (SW) wide-band filters (F090W, F115W, F150W, F200W), 3 long-wavelength (LW) wide-band filters (F277W, F356W, F444W), and 2 LW medium-band filters (F335M, F410M). The JADES observing strategy is described in detail in Eisenstein et al. (2023).
Approximately one observer-frame year later on UT 2024 February 12–18, the Complete NIRCam Grism Redshift Survey (CONGRESS) program (program ID: 3577; PI: Egami) obtained 3-band JWST/NIRCam imaging of GOODS-N, using two SW wide-band filters (F090W, F115W) and one LW wide-band filter (F356W) (Egami et al., 2023). Rieke et al. (2023) provides a detailed description of the data processing procedure for JADES NIRCam observations of GOODS-S. Similar data processing procedures were applied to the JADES GOODS-N NIRCam observations as well as the CONGRESS NIRCam observations (F. Sun et al. in preparation).
@@ 2.2 NIRCam/Grism Spectroscopy
In addition to NIRCam imaging observations, CONGRESS simultaneously obtained 2-hr NIRCam/Grism Wide Field Slitless Spectroscopy (WFSS) observations of GOODS-N using the F356W filter. Previously, the First Reionization Epoch Spectroscopically Complete Observations (FRESCO) program (program ID: 1895; PI: Oesch) obtained 2-hr NIRCam/Grism WFSS observations of GOODS-N using the F444W filter on UT 2023 February 11–13 along with F444W, F182M, and F210M NIRCam imaging observations (Oesch et al., 2023). The spectral resolution of the NIRCam WFSS mode is R 1600 at 2.4–5.0 m.
NIRCam grism data processing was detailed by Sun et al. (2024) following the public codes and calibrations111https://github.com/fengwusun/nircam_grism made available by Sun et al. (2023), including latest updates on tracing and wavelength calibration based on JWST Cycle-1 and 2 data.
@@ 3.2 Photometry
Note. — The F356W magnitude reported for AT 2023adya on TNS is slightly fainter ( 28.270.09) due to the use of an 01 aperture, which failed to capture all of AT 2023adya’s emission.
To verify that the F356W difference image emission at AT 2023adya’s position is real and not a subtraction artifact, we performed aperture photometry at AT 2023adya’s position in the Epoch1 and Epoch2 science images in each filter and measured the brightness variation of the whole galaxy. The two epochs of science image photometry are listed in the top two rows of Table 1. The difference between the F356W Epoch1 and Epoch2 science image photometry yields nearly the exact same value as the photometry measured from the difference image ( 28.040.14 mag), demonstrating that the F356W variability is real and not a subtraction artifact. AT 2023adya and its host were 19% brighter in Epoch1 than Epoch2. We also processed two epochs of imaging data by bootstrapping the NIRCam exposures that cover the position of AT 2023adya, produced multiple mosaicked images at both epochs and performed aperture photometry. Our bootstrapping test validates the F356W variability, which is not an artifact due to certain bad exposures.
@@ 3.3 AT 2023adya’s Position and Host
AT 2023adya was discovered at (18912003, 6223867) in GOODS-N and belongs to host galaxy JADES–GN+189.12004+62.23867. The host resides at 5.274 (Sun et al., 2024, the redshift is updated based on improved calibrations), measured unambiguously from an H emission line in the FRESCO F444W NIRCam/grism spectrum and an [O III] 5008 emission line in the CONGRESS F356W NIRCam/grism spectrum. Sun et al. (2024) measured AT 2024adya’s host to be centered at (18912004, +6223867), which is only 002 (0.10 kpc) away from AT 2023adya’s position. We re-measured the host centroid using the Epoch2 F356W image since Sun et al. (2024) measured the centroid using the AT 2023adya-contaminated images, but we find a negligible difference in the Epoch2 host centroid. Hence, we use the measurement from Sun et al. (2024) as the host centroid. The angular separation between AT 2023adya and its host’s center is 1/3 of a native LW pixel, and thus AT 2023adya can be considered coincident with the host’s center within the uncertainty, as shown in Figure 1.
Although there is 9-band JWST/NIRCam imaging from Epoch1, we do not perform host galaxy SED fitting because those images are contaminated with AT 2023adya, and thus the derived parameters from the SED fit may be unreliable. Epoch2 has only three JWST/NIRCam filters which is insufficient to produce a reliable SED fit.
@@ 4.1.1 SN IIP
At SN 2009kf’s redshift of 0.182, the effective Galaxy Evolution Explorer (GALEX) NUV wavelength corresponds to rest-frame 195nm and thus probes a similar rest-frame wavelength as AT 2023adya’s F115W photometric non-detection. F115W’s limiting magnitude of 28.7 mag corresonds to MNUV 17.8 mag at 5.274, and if AT 2023adya is similar to SN 2009kf, it would have peaked around MNUV 21.5 mag, well above the detection limit. In other words, if AT 2023adya is SN 2009kf-like, it probably would have been detected in F115W. However, we cannot rule out the possibility of dust-extinction causing the NUV non-detection. We are unfortunately unable to provide constraints on potential dust extinction because the 9-band Epoch1 host SED is contaminated with AT 2023adya and thus provides unreliable host characteristics.
@@ 4.2 Active Galactic Nucleus
Optical variability is a common selection technique to identify AGN (Klesman & Sarajedini, 2007; Trevese et al., 2008; Villforth et al., 2010; De Cicco et al., 2015; García-González et al., 2015; Yuk et al., 2022), as many Type 1 AGN exhibit variability on timescales ranging from hours to years (Ulrich et al., 1997; Vanden Berk et al., 2004; Sesar et al., 2007; Hickox et al., 2014). Type 1 AGN exhibit a broad H emission line ( 1000 km/s) that cannot be attributed to normal star-forming galaxy activity. Under the widely-accepted assumption that this broad H emission is connected to AGN activity, the AGN broad line region (BLR) generates the broad H emission.
@@
Theoretical models presented in Hopkins et al. (2006), Somerville et al. (2008), and Hickox et al. (2014) indicate that mergers are a primary mechanism for driving AGN activity. AT 2023adya’s host is a member of a high-confidence merger pair (80% close pair probability; Puskás et al. 2025), residing only 5.2 kpc from JADES-GN-189.11960+62.23855 ( 5.274; Sun et al. 2024), and is also part of overdensity JADES–GN–OD–5.269 (Helton et al., 2024). This suggests that AT 2023adya could result from merger-induced AGN activity.
@@
López-Navas et al. (2023) analyzed the Zwicky Transient Facility light curves of 50% of the AGN from Sloan Digital Sky Survey (SDSS) data release 16 and found that Type 2 AGN show negligible optical variations, and the Type 2 AGN that do show optical variations are likely misclassified Type 1 AGN or changing-look/changing-state AGN. Barth et al. (2014) searched for optical g-band variability among the luminous Type 2 AGN in SDSS Stripe82 and similarly found that the majority of Type 2 AGN do not exhibit optical variability, and most or all of the variable AGN in the Type 2 sample are likely Type 1 contaminants. The lack of a broad H component in AT 2023adya’s host spectrum during Epoch1 means either (1) our view of the AGN BLR is blocked by dust (i.e., Type 2 AGN), or (2) if we have a direct view toward the galaxy center, the galaxy nucleus is not active. In either case, we would not expect to see 19% change in brightness over 2 rest-frame months, suggesting that AT 2023adya is not likely to arise from AGN activity. Although it is difficult to exclude the existence of a faint broad-line component based on the limited-depth NIRCam/Grism spectrum, a 19% brightness change would have made such a component visible in the spectrum.

### v2

@@ Abstract
The James Webb Space Telescope (JWST) discovered 79 transients out to 4.8 through the JADES Transient Survey (JTS), but the JTS did not find any 5 transients. We present the first photometric evidence of a 5 transient/variable source with JWST. The source, AT 2023adya, resides in a 5.274 galaxy in GOODS-N, which dimmed from 26.050.02 mag to 26.240.02 mag in the rest-frame optical over approximately two rest-frame months, producing a clear residual signal in the difference image ( 28.010.17 mag; SNvar 6.09) at the galaxy center. Shorter-wavelength bands (F090W/F115W) show no rest-frame ultraviolet brightness change. Based on its rest-frame V-band absolute magnitude (MV 18.48 mag), AT 2023adya could be any core-collapse supernova (SN) subtype or an SN Ia. However, due to low SN Ia rates at high redshift, the SN Ia scenario is unlikely. Alternatively, AT 2023adya may be a variable active galactic nucleus (AGN). The NIRCam/Grism spectrum shows no broad H emission line (FWHM 13026 km s-1), but we cannot exclude the existence of a faint broad line and therefore we cannot exclude the AGN scenario. AT 2023adya is unlikely to be a tidal disruption event (TDE) because the TDE models matching the observed brightness changes have low event rates. Although it is not possible to determine AT 2023adya’s nature based on the two-epoch single-band photometry alone, this discovery pushes the transient/variable science frontier past 5 and towards the epoch of reionization.
@@ 1 Introduction
Building a sample of high-redshift SNe Ia can provide useful constraints on SN Ia systematics (Riess & Livio, 2006), including whether they are truly similar to low-redshift SNe Ia and can act as standard candles (Pierel et al., 2024a, 2025). This has important implications for cosmology constraints achieved at higher redshifts, such as the possibility of an evolving dark energy equation of state parameter, w (Adame et al., 2025). CCSNe, on the other hand, closely trace the instantaneous star formation rate and thus provide an independent view of the cosmic star formation rate density. Observed CCSN rates can be compared to those expected from the cosmic star formation rate density to provide constraints on missing CCSN fractions, the progenitor mass range of CCSNe, and the potential scenario of an evolving initial mass function (IMF; Dahlen et al. 2012; Strolger et al. 2015). Additionally, there is evidence that CCSN explosion energies evolve with redshift (Coulter et al., 2025; Moriya et al., 2025). Building a larger sample of high-redshift CCSNe will generate better constraints on how high-redshift CCSNe differ from local CCSNe.
@@
The JWST Advanced Deep Extragalactic Survey (JADES) program obtained two epochs of deep (30 mag) JWST near-infrared camera (NIRCam) images of the Great Observatories Origins Deep Survey South (GOODS-S) field, separated by one observer-frame year (Eisenstein et al., 2023). The JADES Transient Survey (JTS) discovered 38 SNe at 2, and the highest-redshift SN candidate in the JTS sample is associated with a 4.82 host (DeCoursey et al., 2025). This sample includes an SN Ia at 2.9 (Pierel et al., 2024a), a Type Ic broad line (Ic-BL) SN at 2.83 (Siebert et al., 2024), and a likely Type IIP SN at 3.61 (Coulter et al., 2025). Moriya et al. (2025) explores the properties of the robustly-classified Type II SNe in the JTS sample, three of which are at 2. Additionally, within the COSMOS-Web SN survey, an SN Ia has been discovered at 2.15 (Pierel et al., 2025). There have also been multiple serendipitous SN discoveries with JWST (e.g. Yan et al., 2023), including multiply-lensed SNe Ia at 1.78 (Frye et al., 2024; Pierel et al., 2024b; Pascale et al., 2025) and 1.95 (Pierel et al., 2024c).
@@
We conducted a systematic transient and variable search in the GOODS-North field with two sets of overlapping F090W, F115W, and F356W JWST/NIRCam imaging, separated by one observer-frame year (DeCoursey et al., 2024). This search yielded a sample of 39 transient and variable sources, one of which is AT 2023adya. AT 2023adya is associated with a 5.274 host galaxy (Sun et al., 2024), making it the highest-redshift transient/variable source showing a photometric brightness change discovered thus far with JWST. We explore the types of transient or variable sources that may explain AT 2023adya, which include CCSNe, SNe Ia, variable AGN, and tidal disruption events (TDEs).
@@
In Section 2, we describe the NIRCam observations covering AT 2023adya. Section 3 explains how we discovered AT 2023adya and presents the photometric measurements of AT 2023adya. We also describe AT 2023adya’s host in Section 3. In Section 4, we step through the possible sources of AT 2023adya’s transient or variable nature, and we summarize our findings in Section 5. Throughout this paper, we express magnitudes using the AB system (Oke & Gunn, 1983) and adopt a flat CDM cosmology with the following parameters: H0 70 km s-1 Mpc-1, 0.7, and 0.3.
@@ 2.1 NIRCam Imaging
The JADES program (program ID: 1181; PI: Eisenstein) obtained 9-band JWST/NIRCam imaging of GOODS-N on UT 2023 February 3, using 4 short-wavelength (SW) wide-band filters (F090W, F115W, F150W, F200W), 3 long-wavelength (LW) wide-band filters (F277W, F356W, F444W), and 2 LW medium-band filters (F335M, F410M). The JADES observing strategy is described in detail in Eisenstein et al. (2023), and the JADES images are publicly available on MAST/high-level science product (HLSP; JADES Team 2024), where AT 2023adya’s host has ID 1026307.
Approximately one observer-frame year later on UT 2024 February 12–18, the Complete NIRCam Grism Redshift Survey (CONGRESS) program (program ID: 3577; PI: Egami) obtained 3-band JWST/NIRCam imaging of GOODS-N, using two SW wide-band filters (F090W, F115W) and one LW wide-band filter (F356W) (Egami et al., 2023). The CONGRESS data is publicly available on MAST (CONGRESS Team, 2025). Rieke et al. (2023) provides a detailed description of the data processing procedure for JADES NIRCam observations of GOODS-S. Similar data processing procedures were applied to the JADES GOODS-N NIRCam observations as well as the CONGRESS NIRCam observations (F. Sun et al. in preparation).
@@ 2.2 NIRCam/Grism Spectroscopy
In addition to NIRCam imaging observations, CONGRESS simultaneously obtained 2-hr NIRCam/Grism Wide Field Slitless Spectroscopy (WFSS) observations of GOODS-N using the F356W filter. Previously, the First Reionization Epoch Spectroscopically Complete Observations (FRESCO) program (program ID: 1895; PI: Oesch) obtained 2-hr NIRCam/Grism WFSS observations of GOODS-N using the F444W filter on UT 2023 February 11–13 along with F444W, F182M, and F210M NIRCam imaging observations (Oesch et al., 2023; FRESCO Team, 2023). The spectral resolution of the NIRCam WFSS mode is R 1600 at 2.4–5.0 m.
NIRCam grism data processing was detailed by Sun et al. (2024) following the public codes and calibrations111https://zenodo.org/records/14052875 (catalog doi:10.5281/zenodo.14052875) (Sun, 2024) made available by Sun et al. (2023), including latest updates on tracing and wavelength calibration based on JWST Cycle-1 and 2 data.
@@ 3.2 Photometry
Note. — The F356W difference image magnitude reported for AT 2023adya on TNS is slightly fainter ( 28.270.09) due to the use of an 01 aperture, which failed to capture all of AT 2023adya’s emission.
To verify that the F356W difference image emission at AT 2023adya’s position is real and not a subtraction artifact, we performed aperture photometry at AT 2023adya’s position in the Epoch1 and Epoch2 science images in each filter and measured the brightness variation of the whole galaxy. The two epochs of science image photometry are listed in the top two rows of Table 1. The difference between the F356W Epoch1 and Epoch2 science image photometry yields nearly the exact same value as the photometry measured from the difference image ( 28.040.14 mag), demonstrating that the F356W variability is real and not a subtraction artifact. AT 2023adya and its host were 19% brighter in Epoch1 relative to Epoch2.
As an additional test to ensure the F356W brightness change is real, we randomly drew individual exposures covering AT 2023adya’s position from the JADES and CONGRESS datasets to produce bootstrapped Epoch1 and Epoch2 mosaics. There were 8 (7) total exposures in the Epoch1 JADES (Epoch2 CONGRESS) data covering AT 2023adya. So, to create the bootstrapped mosaics, we randomly chose 7 (6) exposures and removed duplicates, resulting in 5–7 (4–6) exposures used to generate the bootstrapped Epoch1 JADES (Epoch2 CONGRESS) mosaics. We then performed aperture photometry on each bootstrapped mosaic and compared the results to the galaxy fluxes measured from the full JADES and CONGRESS mosaics (rows 1–2 of Table 1). We performed this exercise six times and found that the measured fluxes from the bootstrapped Epoch1 and Epoch2 mosaics were always consistent with the JADES and CONGRESS galaxy fluxes within 1, respectively. In other words, our bootstrapping test validates that the F356W brightness change is not an artifact due to certain bad exposures.
@@ 3.3 AT 2023adya’s Position and Host
AT 2023adya was discovered at (18912003, 6223867) in GOODS-N and belongs to host galaxy JADES–GN+189.12004+62.23867. The host resides at 5.274 (Sun et al., 2024, the redshift is updated based on improved calibrations), measured unambiguously from an H emission line in the FRESCO F444W NIRCam/Grism spectrum and an [O III] 5008 emission line in the CONGRESS F356W NIRCam/Grism spectrum. Sun et al. (2024) measured AT 2024adya’s host to be centered at (18912004, +6223867), which is only 002 (0.10 kpc) away from AT 2023adya’s position. We re-measured the host centroid using the Epoch2 F356W image since Sun et al. (2024) measured the centroid using the AT 2023adya-contaminated images (i.e., Epoch1 images), but we find a negligible difference in the Epoch2 host centroid. Hence, we use the measurement from Sun et al. (2024) as the host centroid. The angular separation between AT 2023adya and its host’s center is 1/3 of a native LW pixel, and thus AT 2023adya can be considered coincident with the host’s center within the uncertainty, as shown in Figure 1.
Although the 9-band JADES JWST/NIRCam imaging from Epoch1 is contaminated with AT 2023adya, we performed host galaxy spectral energy distrubtion (SED) fitting with a modified version of Prospector (Johnson et al., 2021; Lyu et al., 2022) to estimate the host properties and attempt to identify whether or not an AGN is present (Figure 2). The contamination from AT 2023adya is unlikely to significantly skew the fit because it represents 20% of the F356W flux and does not affect the F090W or F115W flux. The contamination from AT 2023adya presumably affects the other LW filters to a similar extent as the 20% contribution to the F356W flux, though we cannot confirm this. We used the cataloged JADES photometry (JADES Team, 2024) and HST Advanced Camera for Surveys (ACS) photometry measured from the Hubble Legacy Fields (HLF) data release v2.5 (HLF Team, 2015) to perform the fit.
Although our photometric coverage is limited in wavelength, the best-fit SED does not rise toward longer wavelength, showing no obvious sign of AGN contribution to the total emission; the best-fit model spectrum arises solely from stellar contributions. Based on the best-fit model, the host has a mass of log(M/M⊙) 9.13, a metallicity of log(Z/Z⊙) -1.82, and a V-band stellar dust attenuation of AV 0.2. The computed star formation rate is 7.8 M⊙ year-1 assuming a delayed tau star formation history with a stellar age at 0.27 Gyr and a characteristic timescale of 0.10 Gyr.
@@ 4.1.1 SN IIP
At SN 2009kf’s redshift of 0.182, the effective Galaxy Evolution Explorer (GALEX) NUV wavelength corresponds to rest-frame 195nm and thus probes a similar rest-frame wavelength as AT 2023adya’s F115W photometric non-detection. F115W’s limiting magnitude of 28.7 mag corresonds to MNUV 17.8 mag at 5.274, and if AT 2023adya is similar to SN 2009kf, it would have peaked around MNUV 21.5 mag, well above the detection limit. In other words, if AT 2023adya is SN 2009kf-like, it probably would have been detected in F115W. However, we cannot rule out the possibility of dust-extinction causing the NUV non-detection.
@@ 4.2 Active Galactic Nucleus
4.2.1 Type 1 AGN
Optical variability is a common selection technique to identify AGN (Klesman & Sarajedini, 2007; Trevese et al., 2008; Villforth et al., 2010; De Cicco et al., 2015; Garc´ıa-González et al., 2015; Yuk et al., 2022), as many Type 1 AGN exhibit variability on timescales ranging from hours to years (Ulrich et al., 1997; Vanden Berk et al., 2004; Sesar et al., 2007; Hickox et al., 2014). Type 1 AGN exhibit a broad H emission line ( 1000 km/s) that cannot be attributed to normal star-forming galaxy activity. Under the widely-accepted assumption that this broad H emission is connected to AGN activity, the AGN broad line region (BLR) generates the broad H emission.
@@ 4.2.1 Type 1 AGN
It is difficult to exclude the existence of a faint broad-line component based on the limited-depth NIRCam/Grism spectrum. In fact, Zhang et al. (2025) mentions two sources that exhibit no broad features in NIRCam/Grism spectra but are reported to show broad line features in NIRSpec medium-resolution grating data (Maiolino et al., 2024). Zhang et al. (2025) suggests that broad line searches in NIRCam/Grism data are effective down to 26 mag, and the Epoch1 F444W magnitude of AT 2023adya’s host is slightly fainter than that limit (26.4 mag).
To test whether or not we would expect to detect broad H in the NIRCam/Grism data for various Type 1 AGN configurations, we roughly estimated the expected H line flux based on the observed F356W flux density using Equation 1 from Greene & Ho (2005). At 5.274, F356W corresponds to rest-frame 5674 Å, which we approximated as 5100 Å. We tested two different scenarios: (1) AT 2023adya is a Type 1 AGN whose continuum is AGN-dominated, meaning that the observed Epoch1 F356W flux density ( 138 nJy) is entirely from the AGN, and (2) AT 2023adya is a Type 1 AGN whose continuum is dominated by star formation, meaning that the AGN only contributes to the observed change in F356W brightness ( 23 nJy).
In the first scenario (pure Type 1 AGN), the AGN flux density of 138 nJy corresponds to an expected H line flux of 4 10-18 erg s-1 cm-2, which would constitute 60% of the observed H line flux. Given this significant contribution to the total line flux, we would expect to see a broad component in the H line if AT 2023adya were a pure Type 1 AGN. Therefore, our lack of broad H detection means that it is unlikely that AT 2023adya is a Type 1 AGN whose continuum is dominated by the AGN. Conversely, if AT 2023adya is a Type 1 AGN whose continuum is dominated by star formation, the 23 nJy from the AGN would constitute an H line flux of only 6.5 10-19 erg s-1 cm-2, which is roughly ten times less than the measured H line flux. In this case, it is unlikely that we would detect the broad component of H, especially if the line is significantly broadened, given NIRCam/Grism’s limited sensitivity. Therefore, we cannot exclude the possibility that AT 2023adya is a star formation-dominated Type 1 AGN.
4.2.2 Type 2 AGN
According to the simple AGN Unified Model (Antonucci, 1993), the central engines and broad line regions of Type 2 AGN are obscured by toroidal dust distributions. These dusty tori highly suppress optical variations, so we expect Type 2 AGN to exhibit little to no optical variability. Even if we could see the AGN continuum through scattered light, the non-variable stellar emission from the host would dominate and dilute the variability of the scattered AGN light. Barth et al. (2014) validated these predictions by searching for optical g-band variability among the luminous Type 2 AGN in the Sloan Digital Sky Survey (SDSS) Stripe82 field. They found that 90% of the Type 2 AGN did not show optical variability, and of the 10% that did show variability, most if not all of them are misclassified Type 1 AGN. Notably, however, the misclassified Type 1 AGN in Barth et al. (2014) exhibited stronger variability than AT 2023adya.
There is a unique subsample of Type 2 AGN that exhibit similar levels of optical variability as AT 2023adya. These are the so-called “true Type 2” or “naked Type 2” AGN, which are unobscured AGN (i.e., viewed face-on) that intrinsically lack a BLR (Hawkins, 2004). Previous AGN variability surveys such as those conducted by Cartier et al. (2015) and Sánchez et al. (2017) have reported true Type 2 AGN candidates. Furthermore, Barth et al. (2014) were unable to exclude the true Type 2 AGN scenario for a small subset of their variable Type 2 AGN sample, although they disfavor this explanation. We are also unable to exclude the possibility that AT 2023adya is a true Type 2 AGN, but we do not favor this possibility due to the rarity of these objects (Barth et al., 2014).